# Vacancy-Induced Quantum Properties in 2D Silicon Carbide: Atomistic insights from semi-local and hybrid DFT calculations

### v1

@@ Abstract
Two-dimensional (2D) materials have emerged as promising platforms for quantum technologies and optoelectronics, with defects playing a crucial role in their properties. We present a comprehensive density functional theory study of silicon and carbon vacancies in monolayer silicon carbide (1L-SiC), a wide-bandgap 2D semiconductor with potential for room-temperature quantum applications. Using PBE, SCAN, r2SCAN, and HSE06 functionals, we reveal distinct characteristics between Si and C vacancies. Formation energies and charge transition levels show strong functional dependence, with HSE06 consistently predicting higher values and deeper transition levels compared to PBE calculations. Electronic structure analysis demonstrates contrasting behavior: silicon vacancies create highly localized states with strong spin polarization, while carbon vacancies produce more dispersed states with weaker magnetic properties. Vacancy migration studies reveal significantly lower barriers for silicon vacancies compared to carbon vacancies, indicating higher mobility for Si vacancies at moderate temperatures. Optical properties, calculated using PBE-DFPT, show distinct charge-state dependent absorption in the far-infrared region, with positively charged states of both vacancy types demonstrating the strongest response. The complementary characteristics of Si and C vacancies - localized versus dispersed states, different magnetic properties, and distinct optical responses - suggest possibilities for defect engineering in quantum and optoelectronic applications. Our results highlight the critical importance of advanced functionals in accurately describing defect properties and provide a comprehensive framework for understanding vacancy behavior in 2D materials.
I Introduction
The field of two-dimensional (2D) semiconductors has witnessed remarkable progress in recent years, driven by the discovery of novel materials and their intriguing properties. These atomically thin materials, characterized by their unique electronic and optical characteristics, have opened up exciting avenues for various applications, particularly in the realm of quantum information science and optoelectronics. Among the prominent 2D semiconductors, silicon carbide (SiC) has emerged as a promising candidate due to its exceptional quantum properties at room temperature. Defects within SiC, such as silicon vacancies (VSi) and carbon antisites (SiC), play a pivotal role in enabling single-photon emission, a crucial aspect of quantum communications [1, 2, 3]. Through advanced spectroscopic techniques like electron spin resonance (ESR) and photoluminescence (PL), researchers have gained valuable insights into the nature of these defect states and their impact on the material’s electronic and optical properties [4, 5, 6, 7, 8, 9, 10]. While significant advancements have been made in the study of 3D SiC, exploring the two-dimensional form presents an exciting opportunity. Theoretical predictions suggest that local defects in 2D SiC could outperform their 3D counterparts in terms of photon emission efficiency [11, 12, 13]. Furthermore, the existence of Stone-Wales defects, created by rotating a carbon-carbon bond in the SiC lattice, has been proposed, offering superior radiative transitions compared to isolated point defects [14]. Despite the promising theoretical predictions, realizing the potential of 2D SiC in practice remains a significant challenge due to difficulties in synthesizing high-quality 2D SiC and controlling defects at the atomic level [15, 16, 17, 18, 19, 20]. To characterize and study the defects in 2D SiC, advanced techniques with atomic-scale resolution are essential. Scanning tunneling microscopy (STM), high-resolution transmission electron microscopy (HRTEM), and specialized optical characterization methods like photoluminescence (PL) spectroscopy have been instrumental in probing the atomic structure and properties of defects [7, 21, 22, 23, 24, 25, 26]. Substantial progress has been achieved in comprehending the dynamics of defects in 2D SiC via advanced theoretical research, particularly with respect to thermodynamic stability [27, 28, 29]. First-principles calculations based on density functional theory (DFT) have provided valuable insights into the formation energies, electronic structure, and optical properties of various defects in 2D SiC. By comparing the calculated defect formation energies and charge transition levels with experimental data, researchers can identify the most stable defect configurations and predict their behavior under different environmental conditions. One notable theoretical achievement is the prediction of the stability of the divacancy-complex (VSi-VC) in 2D SiC. DFT calculations using hybrid functionals, such as HSE06, have shown that the divacancy-complex exhibits remarkably low formation energy and is thermodynamically favorable compared to isolated vacancies [27]. This finding has important implications for experimentalists aiming to engineer and control defects in freestanding 2D SiC, as the divacancy-complex can serve as a stable and optically active center for quantum applications.

### v2

@@ Abstract
Two-dimensional (2D) materials have emerged as promising platforms for quantum technologies and optoelectronics, with defects playing a crucial role in their properties. We present a comprehensive density functional theory study of silicon and carbon vacancies in monolayer silicon carbide (1L-SiC), a wide-bandgap 2D semiconductor with potential for room temperature quantum applications. Using PBE, SCAN, r2SCAN, and HSE06 functionals, we reveal distinct characteristics between Si and C vacancies. Formation energies and charge transition levels show strong functional dependence, with HSE06 consistently predicting higher values and deeper transition levels compared to PBE calculations. Electronic structure analysis demonstrates contrasting behavior: silicon vacancies create highly localized states with strong spin polarization, while carbon vacancies produce more dispersed states with weaker magnetic properties. Vacancy migration studies reveal significantly lower barriers for silicon vacancies compared to carbon vacancies, indicating higher mobility for Si vacancies at moderate temperatures. Optical properties, calculated using PBE-DFPT, show distinct charge-state dependent absorption in the far-infrared region, with positively charged states of both vacancy types demonstrating the strongest response. The complementary characteristics of Si and C vacancies-localized versus dispersed states, different magnetic properties, and distinct optical responses-suggest possibilities for defect engineering in quantum and optoelectronic applications. Our results highlight the critical importance of advanced functionals in accurately describing defect properties and provide a comprehensive framework for understanding vacancy behavior in 2D materials.
Delaware Energy Institute, University of Delaware, 221 Academy St, Newark, DE 19711, USA
1 Introduction